media%2Ff9c%2Ff9cfb044-eccd-4f3f-af47-1d
question 13 14
Thanks for your help

QUESTIONS AND PROBLEMS 69 Sales Food Beverage Total sales 630,000 140,000 $770,000 Cost of Sales Food Beverages Total costs 252,000 35,000 Gross Profit Controllable Expenses 287,000 $483,000 Salaries and wages Employee benefits Other controllable $173,250 45,045 82,000 expenses Total Controllable Expenses Income before Occupancy $300,295 $182,705 Costs, Interest, Depreciation, and Income Taxes Occupancy Costs Income before Interest 64,000 $118,705 Depreciation, and Income Taxes Interest Depreciation Total Restaurant Profit 10,000 28,500 $38,500 80,205 g. Occupancy costs will increase by $5,000 h. Interest and depreciation will remain the same. Given these anticipated changes, prepare an operating budget for the Downtowner Restaurant for the coming year erring to Question 13, assume that the operating budget you pre- pared for the upcoming year has been adopted. After the first six months of the year, financial records reveal the following a. Food percent rather than by the 10 per- sales have increased by 12 cent anticipated. b. Beverage sales have increased by 5 percent rather than by 6 percent.
CHAPTER 2 THE CONTROL PROCESS c. Food cost percent is 1 percent lower than budgeted, but beverage cost percent is 2 percent higher d. Variable salaries and wages are 14 percent of food sales, rather than the 16 percent anticipated. Assuming that the trends evident in the first six months continue for the rest of the year and that both sales and costs are equally divided between the two halves of the year, prepare a revision of the budget for the second six-month period.
Answer #1

Lets solve this problem in steps

1) First, we need to calculate the required percentages from the current year operating budget as follows:-

Percentage of Food cost over its sales = (Food Cost/Food Sales)*100 = ($252,000/$630,000)*100 = 40%

Percentage of Beverages cost over its sales = (Beverage cost/Beverage Sales)*100 = ($35,000/$140,000)*100 = 25%

Employee benefits % over salaries & wages in current year = Employee Benefits/Salaries & Wages

= $45,045/$173,250 = 0.26 or 26%

2) Operating budget for the Downtowner Restaurant for the coming year is shown as follows:-

Downtowner Restaurant

Operating Budget for Coming Year (Amounts in Dollars)

Sales
Food (630,000*1.10) 693,000
Beverages (140,000*1.06) 148,400
Total Sales (A) 841,400
Cost of Sales
Food (693,000*40%) 277,200
Beverages (148,400*25%) 37,100
Total costs (B) (314,300)
Gross Profit (C = A-B) 527,100
Controllable Expenses
Salaries and Wages [69,300+8,000+(693,000*16%)] 188,180
Employee Benefits (188,180*26%) 48,927
Other Controllable expenses (Bal fig) (312,295-188,180-48,927) 75,188
Total Controllable expenses (300,295+12,000) (D) (312,295)
Income before occupancy costs, Interest, Dep. and income taxes (E = C-D) 214,805
Less: Occupancy Costs (64,000+5,000) (69,000)
Income before Interest, Dep. and income taxes 145,805
Interest 10,000
Depreciation 28,500
Total (38,500)
Restaurant Profit 107,305

Working Notes:-

1) Total controllable expenses increase by 12,000 from 300,295 to 312,295, therefore other controllable expenses will be equal to total controllable expenses minus expected salaries and wages and expected employee benefits.

2) Expected variable salaries and wages is 16% of expected food sales (i.e. 693,000*16%).